# Toric varieties modulo reflections

### v1

@@ Abstract.
Let $W$ be a finite group generated by reflections of a lattice $M$. If a lattice polytope $P\subset M\otimes_{\mathbb{Z}}\mathbb{R}$ is preserved by $W$, then we show that the quotient of the projective toric variety $X_{P}$ by $W$ is isomorphic to the toric variety $X_{P\cap D}$, where $D$ is a fundamental domain for the action of $W$. This answers a question of Horiguchi-Masuda-Shareshian-Song, and recovers results of Blume, of the second author, and of Gui-Hu-Liu.
@@ 1. Introduction
1.1.
The quotient of a toric variety by a finite group need not be toric, even when the group acts by toric automorphisms. For instance, the quotient of $\mathbb{P}^{n}\times\mathbb{P}^{n}$ by the $\mathbb{Z}/2$ action that swaps the factors is not toric when $n\geq 2$ [HMJJ21, Remark 8.2], and the quotient of a 47-torus over $\mathbb{Q}$ by the $\mathbb{Z}/47$ action that cycles its factors is not even rational [S69, Theorem 1] [^1] We thank Laurent Moret-Bailly for writing the MathOverflow post that made us aware of this example.! The existence of these examples is unsuprising: it is almost never possible to make the quotient map torus-equivariant, so there is little reason to expect a toric structure on the quotient.
[…]
@@
for some nonzero $\alpha\in V$. A reflection group in $V$ is a finite subgroup of $\operatorname{GL}(V)$ generated by reflections. The fundamental domain for the action of a finite reflection group $W$ in $V$ is a polyhedral cone $D\subset V$.
@@
[L05, Theorem 6.1.1] If $W\subset\operatorname{GL}(V)$ is a finite reflection group preserving a lattice $M\subset V$, then there is an isomorphism $\mathbb{Z}[D\cap M]\to\mathbb{Z}[M]^{W}$.
The semigroup $D\cap M$ is an example of a saturated affine semigroup, i.e. a finitely generated $\mathbb{Z}_{\geq 0}$-submodule $S\subset M$ such that for all $k\in\mathbb{Z}_{\geq 0}$ and $m\in M$, $km\in S$ implies $m\in S$. Saturated affine semigroup algebras are precisely the the coordinate rings of normal affine toric varieties [CLS11, Theorem 1.3.5], so we may restate Theorem 1.1 as:
@@
[L05, Theorem 6.1.1] If $W\subset\operatorname{GL}(V)$ is a reflection group preserving a lattice $M\subset V$, and $T=\operatorname{Spec}\mathbb{Z}[M]$, then there is an isomorphism of algebraic varieties $T/W\cong\operatorname{Spec}\mathbb{Z}[D\cap M]$.
@@
The main result of this note is a generalization of Lorenz’s theorem.
@@
Equivalently, we may state Theorem 1.2 as:
@@
Projective toric varieties with dense torus $\operatorname{Spec}\mathbb{Z}[M]$ correspond to full-dimensional lattice polytopes $P\subset V$ [CLS11, Chapter 2]. A consequence of Theorem 1.2 is the following statement, conjectured in [HMJJ21, Question 8.1].
@@
Let $W\subset\operatorname{GL}(V)$ be a reflection group preserving a lattice $M\subset V$ and a full-dimensional lattice polytope $P\subset V$, with associated projective toric variety $X_{P}$. The quotient $X_{P}/W$ is isomorphic to the projective toric variety $X_{P\cap D}$.
@@
Special cases of Corollary 1.3 include the well-known isomorphisms $(\mathbb{P}^{1})^{n}/S_{n}\cong\mathbb{P}^{n}$ and $\mathbb{P}^{n-1}/S_{n}\cong\mathbb{P}(1,2,\ldots,n)$, as well as results of Blume [B15, Theorems 3.1, 5.3, 5.9], which state that the projective toric variety associated to a permutohedron of type A, B, or C modulo the corresponding Weyl group is toric. Corollary 1.3 also strengthens the main theorem of the second author [G24], which says that the quotient of the toric variety of a dominant weight polytope by a parbolic subgroup of a Weyl group of appropriate type is homotopy equivalent to a toric variety. Passing to cohomology rings, Corollary 1.3 recovers the main theorem of [HMJJ21], as well as [GHL24, Theorems 1.2] for crystallographic reflection groups. Unlike [HMJJ21, GHL24], we do not produce an explicit isomorphism $H^{*}(X_{P\cap D})\to H^{*}(X_{P})^{W}$.
Future work
In forthcoming work, we will treat the case of arbitrary fans. In particular, we will describe the normal fan of the polytope $P\cap D$ and consider quotients of toric varieties that are neither affine nor projective.
@@ Acknowledgements
We are grateful to Tao Gui for making us aware of [HMJJ21, Question 8.1]. We also thank Chiara Damiolini & Sam Payne for discussing early versions of this project. CC and CS thank the Simons Foundation for support; CC is also supported by NSF grant DMS-2039316.
@@ 2. Finite reflection groups
We review the theory of finite reflection groups, following [H90].
@@
[…]
@@
    $W$ is generated by $\{s_{\alpha}:\alpha\in\Delta\}$ [H90, Theorem 1.5].
@@
[…]
@@
Remark 2.2.
The definition of “root system” used here is that of [H90], but differs slightly from the definition used in some other sources. To assure the reader that this discrepancy causes no problem, we will provide proofs for results from references other than [H90].
@@
Fix a simple system $\Delta$ for $W$, with corresponding fundamental domain $D$. Define a partial order on $V$ by $u\geq v$ if and only if $u-v$ is a non-negative real linear combination of simple roots.
@@
[H90, Lemma 1.12] If $v\in D$, then $v-wv\in\mathbb{Z}_{\geq 0}\left\langle\alpha_{1},\ldots,\alpha_{r}\right\rangle$ for all $w\in W$. In particular, $v\geq wv$.
@@
[BtD85, Proposition VI.2.4(ii)] If $u,v\in D$, then $u\geq v$ if and only if $\operatorname{conv}(Wu)$ contains $v$.
@@
If $v\in\operatorname{conv}(Wu)$, then there are non-negative real numbers $c_{w}$, $w\in W$, such that $\sum_{w}c_{w}=1$ and $v=\sum_{w}c_{w}wu$. By Proposition 2.3, $wu\leq u$, so
@@
Conversely, suppose $v\not\in\operatorname{conv}(Wu)$. Then there is $l\in V$ such that for all $w\in W$, $\left\langle v,l\right\rangle>\left\langle wu,l\right\rangle=\left\langle u,w^{-1}l\right\rangle$. In particular, if $l^{\prime}=w^{-1}l\in D$, then by Proposition 2.3,
@@
[…]
@@
[speyernotes, §15] Let $W$ be a finite reflection group that preserves a lattice $M\subset V$. There is a simple system $\Delta$ for $W$ such that
@@
Morevoer, the action of $W$ preserves $\Lambda$.
@@ 3. Proofs
[…]
@@
  • $W\subset\operatorname{GL}(V)$ is a finite reflection group preserving $M$, with simple system $\Delta=\{\alpha_{1},\ldots,\alpha_{r}\}\subset M$ as in Proposition 2.5,
@@
The decompositions in Section 2.4 imply that
@@
$\mathbb{Z}[\Lambda]^{W}$ $\cong\mathbb{Z}[Z]\otimes_{\mathbb{Z}}\mathbb{Z}[\Lambda_{\circ}]^{W}=\mathbb{Z}[\Lambda]^{W}\cong\mathbb{Z}[Z]\otimes_{\mathbb{Z}}\mathbb{Z}[\chi^{\lambda_{1}},\ldots,\chi^{\lambda_{r}}]^{W},$
@@
[B02, Theorem VI.3.4.1] $\Psi_{\circ}$ is an isomorphism.
Theorem 3.1 implies $\Psi$ is also an isomorphism. Since our proof of Theorem 1.2 already requires analysis of $\Psi$, we provide a proof of Theorem 3.1 in Section 3.3.
@@
for some $(v_{ij})\in\prod_{\begin{subarray}{c}i=1,\ldots,r\\ j=1,\ldots,b_{i}\end{subarray}}W\lambda_{i}$. By Proposition 2.3, $v_{ij}\leq\lambda_{i}$ for all $i$, so
@@
Let $C:=\mathbb{R}_{\geq 0}\left\langle s:s\in S\right\rangle\subset V$. Plainly, $C$ is convex and preserved by $W$. Moreover, $S=C\cap M$ by [CLS11, Theorem 1.3.5].
Let $u\in D\cap S$ and $v\in\operatorname{supp}\Psi(\chi^{v})$. We may write uniquely $u=z+u_{\circ}$ and $v=z+v_{\circ}$, with $v_{\circ}\in\operatorname{supp}\Psi_{\circ}(\chi^{u_{\circ}})$.
We first show that $v\in M$. By Lemma 3.2, $u-v=u_{\circ}-v_{\circ}\in\mathbb{Z}_{\geq 0}\left\langle\alpha_{1},\ldots,\alpha_{r}\right\rangle$. Since $u,\alpha_{1},\ldots,\alpha_{r}\in M$, $v$ must also lie in $M$.
Next, we show that $v\in C$. By Lemma 3.2, $u_{\circ}\geq v_{\circ}$. By Proposition 2.4, this means $v_{\circ}$ is in the convex hull of $Wu_{\circ}$. Consequently,
@@
3.3.
We now move to establish the main results.
@@
We induct on the cardinality of $U$. If #$\#U=1$, the statement is true by Lemma 3.2ii.
Otherwise, if #$\#U>1$, let $v\in U$ be a maximal element with respect to $\leq$. By Lemma 3.2, there is no $v^{\prime}\in U\setminus v$ such that $v\in\operatorname{supp}\Psi_{\circ}(\chi^{v^{\prime}})$; however, $v\in\operatorname{supp}\Psi_{\circ}(\chi^{v})$. Consequently, $\Psi_{\circ}(\chi^{v})$ is not in the span of $\{\Psi_{\circ}(\chi^{u}):u\in U\setminus v\}$. By the induction hypothesis, it follows that $\{\Psi_{\circ}(\chi^{u}):u\in U\}$ is linearly independent. ∎
Proof of Theorem 3.1.
We first show $\Psi_{\circ}$ is injective. If $f\in\ker(\Psi_{\circ})$, then $f$ is contained in a subgroup of the form $\mathbb{Z}\left\langle U\right\rangle:=\mathbb{Z}\left\langle\chi^{u}:u\in U\right\rangle$ with $U\subset D\cap\Lambda_{\circ}$ a finite set. Lemma 3.4 implies that $\Psi_{\circ}$ is injective on $U$, so $f=0$.
@@
We have $n(u)=1$ if and only if $u$ is a minimal element of $D\cap\Lambda_{\circ}$ with respect to $\leq$ (minimal elements exist because for any $u\in D\cap\Lambda_{\circ}$, $\{v\in D\cap\Lambda_{\circ}:v\leq u\}$ is the intersection of the discrete set $\Lambda_{\circ}$ with a compact set by Proposition 2.4, hence finite).
If $n(u)=1$, then Lemma 3.2 implies $\Psi_{\circ}(\chi^{u})=\sum_{v\in Wu}\chi^{v}=\underline{\chi}^{u}$. Otherwise, suppose $n(u)>1$. By the induction hypothesis and Lemma 3.2,
@@
Proof of Theorem 1.2.
By Theorem 3.1, $\Psi_{\circ}$ is an isomorphism. Since $\mathbb{Z}[Z]$ is free over $\mathbb{Z}$, it follows that $\Psi$ is also an isomorphism. It now suffices to show that $\Psi(\mathbb{Z}[D\cap S])=\mathbb{Z}[S]^{W}$.
The containment $\Psi(\mathbb{Z}[D\cap S])\subset\mathbb{Z}[S]^{W}$ holds by Lemma 3.3. For the reverse containment, suppose $f\in\mathbb{Z}[D\cap\Lambda]$ is not in $\mathbb{Z}[D\cap S]$. There there is a maximal $v\in D\cap\Lambda$ such that $v\not\in S$, but $\chi^{v}$ appears in $f$. By Lemma 3.2 and Lemma 3.3, if $\chi^{v^{\prime}}$ appears in $f$ with $v^{\prime}\neq v$, then $v\not\in\operatorname{supp}\Psi(\chi^{v^{\prime}})$; therefore, $v\in\operatorname{supp}\Psi(f)$. It follows that $\Psi(f)$ is not contained in $\mathbb{Z}\left\langle\chi^{s}:s\in S\right\rangle$, hence is also not contained in
@@
Proof of Theorem 1.2.
In light of [CLS11, Theorem 1.3.5], we need only check that the linear functionals $f_{\alpha}=\left\langle\alpha,-\right\rangle$, $\alpha\in\Delta$, defining $D$ can be rescaled to lie in $\operatorname{Hom}(M,\mathbb{Z})$. This is possible because $M\subset\Lambda$, so $\frac{2f_{\alpha}}{\left\langle\alpha,\alpha\right\rangle}$ takes integer values on $M$. ∎
Proof of Corollary 1.3.
Recall from [CLS11, Chapter 2] that a lattice polytope $P\subset V$, gives rise to the projective toric variety $X_{P}:=\operatorname{Proj}\mathbb{Z}[tP]$, where $\mathbb{Z}[tP]$ is the saturated affine semigroup
$S:=\{(t,p):t\in\mathbb{Z}_{\geq 0},p\in tP\}\subset\mathbb{Z}\times M,$
graded by $\deg(t,p)=t$. Since $W$ preserves $M$ and $P$, it acts on $S$, hence on $\mathbb{Z}[S]$, yielding an action on $X_{P}$ by toric automorhpisms. The fundamental domain of the $W$-action on ×$\mathbb{R}\times V$ is ×$\mathbb{R}\times D$, so by Theorem 1.2, there is an isomorphism of graded rings
$\mathbb{Z}[t(D\cap P)]=\mathbb{Z}[(\mathbb{R}\times D)\cap S]\to\mathbb{Z}[S]^{W}.$
@@
For instance, if $u=e_{1}+3e_{2}=4z+2\lambda$, then $\Psi(\chi^{u})=\chi^{4z}\left(\chi^{\lambda}+\chi^{s_{\alpha}(\lambda)}\right)^{2}=\chi^{4z+2\lambda}+2\chi^{4z}+\chi^{4z-2\lambda}$. See Fig. 1. Geometrically, we have that $\mathbb{A}^{2}/{S_{2}}\cong\mathbb{A}^{2}$.
$\alpha$$\lambda$$s_{\alpha}(\lambda)$$z$×$\times 1$×$\times 2$×$\times 1$

### v2

@@ Abstract.
Let $W$ be a finite group generated by reflections of a lattice $M$. If an $M$-rational polytope $P\subset M\otimes_{\mathbb{Z}}\mathbb{R}$ is preserved by $W$, then we show that the quotient of the projective toric variety $X_{P}$ by $W$ is isomorphic to the toric variety $X_{P\cap D}$, where $D$ is a fundamental domain for the action of $W$. This answers a question of Horiguchi-Masuda-Shareshian-Song, and recovers results of Blume, of the second author, and of Gui-Hu-Liu. We also study quotients of real toric varieties, proving that $X_{P}^{\mathbb{R}}/W$ is contractible when $P$ is a permutohedron.
@@ 1. Introduction
$1.$
The quotient of a toric variety by a finite group need not be toric, even when the group acts by toric automorphisms. For instance, the quotient of $\mathbb{P}^{n}\times\mathbb{P}^{n}$ by the $\mathbb{Z}/2$ action that swaps the factors is not toric when $n\geq 2$ [Hor+24, Remark 8.2], and the quotient of a 47-torus over $\mathbb{Q}$ by the $\mathbb{Z}/47$ action that cycles its factors is not even rational [Swa69, Theorem 1] [^1] We thank Laurent Moret-Bailly for writing the MathOverflow post that made us aware of this example.! The existence of these examples is unsurprising: it is almost never possible to make the quotient map torus-equivariant, so there is little reason to expect a toric structure on the quotient.
Nonetheless, Lorenz [Lor05] provides a class of finite group actions on tori with toric quotient. Let $V$ be a finite-dimensional real vector space with a positive definite bilinear form $\left\langle,\right\rangle$. A reflection in $V$ is a linear transformation of the form
@@
for some nonzero $\alpha\in V$. A reflection group in $V$ is a finite subgroup of $\operatorname{GL}(V)$ generated by reflections. A finite reflection group $W$ in $V$ has a fundamental domain that is a polyhedral cone $D\subset V$.
@@
[Lor05, Theorem 6.1.1] If $W\subset\operatorname{GL}(V)$ is a finite reflection group preserving a lattice $M\subset V$, then there is an isomorphism $\mathbb{Z}[D\cap M]\to\mathbb{Z}[M]^{W}$.
The semigroup $D\cap M$ is an example of a saturated affine semigroup, i.e. a finitely generated $\mathbb{Z}_{\geq 0}$-submodule $S\subset M$ such that for all $k\in\mathbb{Z}_{\geq 0}$ and $m\in M$, $km\in S$ implies $m\in S$. Saturated affine semigroup algebras are precisely the coordinate rings of normal affine toric varieties [CLS11, Theorem 1.3.5], so we may restate Theorem˜1.1 as:
@@
[Lor05, Theorem 6.1.1] If $W\subset\operatorname{GL}(V)$ is a reflection group preserving a lattice $M\subset V$, and $T=\operatorname{Spec}\mathbb{Z}[M]$, then there is an isomorphism of algebraic varieties $T/W\cong\operatorname{Spec}\mathbb{Z}[D\cap M]$.
@@
Our first result (illustrated in Example˜3.5) is a generalization of Lorenz’s theorem.
@@
Equivalently, we may state Theorem˜1.2 as:
@@
Projective toric varieties with dense torus $\operatorname{Spec}\mathbb{Z}[M]$ are given by full-dimensional $M$-rational polytopes $P\subset V$[CLS11, Chapter 2]. [^2] An $M$-rational polytope is a polytope with all vertices in a fractional multiple of $M$. We use rational polytopes instead of lattice polytopes, because the latter class is not preserved under intersection with $D$. A consequence of Theorem˜1.2 is the following statement, conjectured in [Hor+24, Question 8.3].
@@
Let $W\subset\operatorname{GL}(V)$ be a reflection group preserving a lattice $M\subset V$ and a full-dimensional $M$-rational polytope $P\subset V$, with associated projective toric variety $X_{P}$. Then the quotient $X_{P}/W$ is isomorphic to the projective toric variety $X_{P\cap D}$.
The quotients we refer to in Theorems˜1.1, 1.2 and 1.3 are “geometric quotients” in the sense of Geometric Invariant Theory (GIT), which always exist for the action of a finite group on a quasi-projective variety [Dol03, Example 6.1]. Over the complex numbers, the analytic topology of $X/W$ is the quotient topology of the $W$-action on the analytic topology of $X$ by [Nee85, Remark 1.7].
Horiguchi, Masuda, Shareshian, and Song were motivated to conjecture Corollary˜1.3 after establishing an isomorphism of rational cohomology rings $H^{*}(X_{P}/W;\mathbb{Q})\cong H^{*}(X_{P\cap D};\mathbb{Q})$ in the case when $W$ is a parabolic subgroup of an irreducible reflection group $W^{\prime}$ of classical type, and $P$ is a $W^{\prime}$-permutohedron (i.e. the vertices of $P$ are $\{w(u):w\in W^{\prime}\}$, with $u\in M$ a point on which $W^{\prime}$ acts freely) [Hor+24, Theorem 1.1]. Gui-Hu-Liu [GHL25, Theorem 1.2] and Song [Son22, Theorem 1.1] subsequently established similar cohomological isomorphisms when $P$ is a permutohedron of any Lie type, and when $P$ is two-dimensional, respectively. Meanwhile, Gong [Gon24, Theorems 1.1, 1.4] obtained a topological result, proving that $X_{P}/W$ and $X_{P\cap D}$ are homotopy equivalent when $P$ is a $W^{\prime}$-permutohedron and $W\subset W^{\prime}$ is a parabolic subgroup. All of these statements are recovered by Corollary˜1.3.
@@
Further special cases of Corollary˜1.3 include the well-known isomorphisms $(\mathbb{P}^{1})^{n}/S_{n}\cong\mathbb{P}^{n}$ and $\mathbb{P}^{n-1}/S_{n}\cong\mathbb{P}(1,2,\ldots,n)$, as well as results of Blume [Blu15, Theorems 3.1, 5.3, 5.9], which state that the projective toric variety associated to a permutohedron of type A, B, or C modulo the corresponding Weyl group is toric.
1.3.
We next study the topological quotient $X_{P}^{\mathbb{R}}/W$ of the real points $X_{P}^{\mathbb{R}}$ of $X_{P}$. The algebraic results described above are of little avail in this setting, as the GIT quotient of a variety by a finite group may not coincide with the topological quotient when working over a field that is not algebraically closed.
Emulating [Son22], we first study the surface case.
Theorem 1.4.
If $\dim P=2$ and $W$ is the Weyl group of an irreducible crystallographic root system of rank 2, then $X_{P}^{\mathbb{R}}/W$ is homotopy equivalent to $\bigvee_{\ell-1}S^{1}$, where $\ell$ is the number of vertices of $P$ that lie in $D$.
We then consider the setting of [GHL25], in which $P$ is a $W$-permutohedron. The topology of real permutohedral varieties is well-studied and combinatorially intricate (see, e.g. [Hen12, CCK19, CY23]). In contrast, we show:
Theorem 1.5.
If $W$ is the Weyl group of an irreducible crystallographic root system, $M$ is its root lattice, and $P$ is a $W$-permutohedron, then $X_{P}^{\mathbb{R}}/W$ is contractible.
An immediate consequence of Theorem˜1.5 is:
Corollary 1.6.
In the notation of Theorem˜1.5, the trivial representation of $W$ is not a subrepresentation of $H^{i}(X_{P}^{\mathbb{R}};\mathbb{C})$ for $i>0$.
Corollary˜1.6 follows in types A and B from works of [Hen12] and [CCK19], respectively, on the $\mathbb{Q}[W]$-module structure of $H^{*}(X_{P}^{\mathbb{R}};\mathbb{Q})$. (We remark that these module structures also imply that quotients of $X_{P}^{\mathbb{R}}$ by parabolic subgroups of $W$ may have non-trivial topology. Describing these quotients remains an interesting problem.)
Motivated by Theorems 1.4 and 1.5, we make the following conjecture on the topology of $X_{P}^{\mathbb{R}}/W$.
Conjecture 1.7.
Suppose $W$ is the Weyl group of an irreducible crystallographic root system that spans $V$, and $M$ is its root lattice. Then $X_{P}^{\mathbb{R}}/W$ is homotopy equivalent to a wedge of spheres, and has Euler characteristic $\chi(X_{P}^{\mathbb{R}}/W)=2$-$\ell$, where $\ell$ is the number of vertices of $P$ that lie in $D$.
Our proofs of Theorems 1.2, 1.2, and Corollary˜1.3, uniform with respect to $W$, $M$, and $P$, stand starkly against the ad-hoc approach we employ to prove Theorems 1.4 and 1.5. We hope that ˜1.7 will inspire the development of a more conceptual approach to the latter results.
Organization
In Section˜2, we review finite reflection groups, then establish Theorems˜1.2, 1.2 and 1.3 in Section˜3. In the remainder of the paper, we turn our attention to real points. We review topological models for $X_{P}^{\mathbb{R}}$ and $X_{P}^{\mathbb{R}}/W$, then prove Theorem˜1.4 in Section˜4. Finally, in Section˜5, we study quotients of real permutohedral varieties and prove Theorem˜1.5.
@@ Acknowledgements
We are grateful to Gui Tao for making us aware of [Hor+24, Question 8.1], and to Hiraku Abe & Haozhi Zeng for comments that improved the exposition. We also thank Chiara Damiolini & Sam Payne for discussing early versions of this project. TG is indebted to Matthias Franz for much help, and to Suyoung Choi, Ajneet Dhillon and Nicole Lemire for discussions. CC and CS thank the Simons Foundation for support; CC is also supported by NSF grant DMS-2039316.
@@ 2. Finite reflection groups
We review the theory of finite reflection groups, following [Hum90].
@@
Let $V$ be a real vector space equipped with a positive definite symmetric bilinear form $\left\langle,\right\rangle$. For $R=\mathbb{Z}_{\geq 0},\mathbb{Z},\mathbb{R}$, we write $R\left\langle v_{1},\ldots,v_{k}\right\rangle$ for the $R$-span of vectors $v_{1},\ldots,v_{k}$ (the use of angle brackets for both span and inner product should cause no confusion). The reflection by a nonzero $\alpha\in V$ is the linear transformation given by
@@
    $W$ is generated by $\{s_{\alpha}:\alpha\in\Delta\}$ [Hum90, Theorem 1.5].
@@
    $D=\{v\in V:\left\langle v,\alpha\right\rangle\geq 0,\;\alpha\in\Delta\}$ is a fundamental domain for the action of $W$ on $V$ (i.e. each $W$-orbit intersects $D$ in exactly one point) [Hum90, Theorem 1.12].
@@
Remark $2.$
The definition of “root system” used here is that of [Hum90], but differs slightly from the definition used in some other sources. To assure the reader that this discrepancy causes no problem, we will provide proofs for results from references other than [Hum90].
@@
Fix a simple system $\Delta$ for $W$, with corresponding fundamental domain $D$. Define a partial order on $V$ by $u\geq v$ if and only if uvu-v is a non-negative real linear combination of simple roots.
@@
[Hum90, Lemma 1.12] If $v\in D$, then $v-wv\in\mathbb{R}_{\geq 0}\left\langle\alpha_{1},\ldots,\alpha_{r}\right\rangle$ for all $w\in W$. In particular, $v\geq wv$.
@@
[BD03, Proposition VI.2.4(ii)] If $u,v\in D$, then $u\geq v$ if and only if $\operatorname{conv}(Wu)$ contains $v$.
@@
If $v\in\operatorname{conv}(Wu)$, then there are non-negative real numbers $c_{w}$, $w\in W$, such that $\sum_{w}c_{w}=1$ and $v=\sum_{w}c_{w}wu$. By Proposition˜2.3, $wu\leq u$, so
@@
Conversely, suppose $v\not\in\operatorname{conv}(Wu)$. Then there is $l\in V$ such that for all $w\in W$, $\left\langle v,l\right\rangle>\left\langle wu,l\right\rangle=\left\langle u,w^{-1}l\right\rangle$. In particular, if $l^{\prime}=w^{-1}l\in D$, then by Proposition˜2.3,
@@
$\Lambda=\{v\in V:2\frac{\left\langle v,\alpha_{i}\right\rangle}{\left\langle\alpha_{i},\alpha_{i}\right\rangle}\in\mathbb{Z},\,1\leq i\leq r\}.$
@@
[Spe, §15] Let $W$ be a finite reflection group that preserves a lattice $M\subset V$. There is a simple system $\Delta$ for $W$ such that
@@
Moreover, the action of $W$ preserves $\Lambda$.
@@
  • $V$ is a real vector space with positive definite symmetric bilinear form $\left\langle,\right\rangle$,
@@
  • $W\subset\operatorname{GL}(V)$ is a finite reflection group preserving $M$, with simple system $\Delta=\{\alpha_{1},\ldots,\alpha_{r}\}\subset M$ as in Proposition˜2.5,
@@
The decompositions in Section˜2.4 imply that
@@
$\mathbb{Z}[\Lambda]^{W}$ $\cong\mathbb{Z}[Z]\otimes_{\mathbb{Z}}\mathbb{Z}[\Lambda_{\circ}]^{W}=\mathbb{Z}[Z]\otimes_{\mathbb{Z}}\mathbb{Z}[\chi^{\lambda_{1}},\ldots,\chi^{\lambda_{r}}]^{W},$
@@
[Bou02, Theorem VI.3.4.1] $\Psi_{\circ}$ is an isomorphism.
Theorem˜3.1 implies $\Psi$ is also an isomorphism. Since our proof of Theorem˜1.2 already requires analysis of $\Psi$, we provide a proof of Theorem˜3.1 in Section˜3.3.
@@
for some $(v_{ij})\in\prod_{\begin{subarray}{c}i=1,\ldots,r\\ j=1,\ldots,b_{i}\end{subarray}}W\lambda_{i}$. By Proposition˜2.3, $v_{ij}\leq\lambda_{i}$ for all $i$, so
@@
Let $C:=\mathbb{R}_{\geq 0}\left\langle s:s\in S\right\rangle\subset V$. Plainly, CC is convex and preserved by $W$. Moreover, $S=C\cap M$ by [CLS11, Theorem 1.3.5].
Let $u\in D\cap S$ and $v\in\operatorname{supp}\Psi(\chi^{u})$. We may write uniquely $u=z+u_{\circ}$ and $v=z+v_{\circ}$, with $v_{\circ}\in\operatorname{supp}\Psi_{\circ}(\chi^{u_{\circ}})$.
We first show that $v\in M$. By Lemma˜3.2, $u-v=u_{\circ}-v_{\circ}\in\mathbb{Z}_{\geq 0}\left\langle\alpha_{1},\ldots,\alpha_{r}\right\rangle$. Since $u,\alpha_{1},\ldots,\alpha_{r}\in M$, $v$ must also lie in $M$.
Next, we show that $v\in C$. By Lemma˜3.2, $u_{\circ}\geq v_{\circ}$. By Proposition˜2.4, this means $v_{\circ}$ is in the convex hull of $Wu_{\circ}$. Consequently,
@@
$3.$
We now move to establish Theorems˜1.2 and 1.2, and Corollary˜1.3.
@@
We induct on the cardinality of $U$. If #U=1\#U=1, the statement is true by Lemma˜3.2Item˜ii.
Otherwise, if #U>1\#U>1, let $v\in U$ be a maximal element with respect to $\leq$. By Lemma˜3.2, there is no $v^{\prime}\in U\setminus v$ such that $v\in\operatorname{supp}\Psi_{\circ}(\chi^{v^{\prime}})$; however, $v\in\operatorname{supp}\Psi_{\circ}(\chi^{v})$. Consequently, $\Psi_{\circ}(\chi^{v})$ is not in the span of $\{\Psi_{\circ}(\chi^{u}):u\in U\setminus v\}$. By the induction hypothesis, it follows that $\{\Psi_{\circ}(\chi^{u}):u\in U\}$ is linearly independent. ∎
Proof of Theorem˜3.1.
We first show $\Psi_{\circ}$ is injective. If $f\in\ker(\Psi_{\circ})$, then $f$ is contained in a subgroup of the form $\mathbb{Z}\left\langle U\right\rangle:=\mathbb{Z}\left\langle\chi^{u}:u\in U\right\rangle$ with $U\subset D\cap\Lambda_{\circ}$ a finite set. Lemma˜3.4 implies that $\Psi_{\circ}$ is injective on $U$, so $f=0$.
@@
We have $n(u)=1$ if and only if $u$ is a minimal element of $D\cap\Lambda_{\circ}$ with respect to $\leq$ (minimal elements exist because for any $u\in D\cap\Lambda_{\circ}$, $\{v\in D\cap\Lambda_{\circ}:v\leq u\}$ is the intersection of the discrete set $\Lambda_{\circ}$ with a compact set by Proposition˜2.4, hence finite).
If $n(u)=1$, then Lemma˜3.2 implies $\Psi_{\circ}(\chi^{u})=\sum_{v\in Wu}\chi^{v}=\underline{\chi}^{u}$. Otherwise, suppose $n(u)>1$. By the induction hypothesis and Lemma˜3.2,
@@
Proof of Theorem˜1.2.
By Theorem˜3.1, $\Psi_{\circ}$ is an isomorphism. Since $\mathbb{Z}[Z]$ is free over $\mathbb{Z}$, it follows that $\Psi$ is also an isomorphism. It now suffices to show that $\Psi(\mathbb{Z}[D\cap S])=\mathbb{Z}[S]^{W}$.
The containment $\Psi(\mathbb{Z}[D\cap S])\subset\mathbb{Z}[S]^{W}$ holds by Lemma˜3.3. For the reverse containment, suppose $f\in\mathbb{Z}[D\cap\Lambda]$ is not in $\mathbb{Z}[D\cap S]$. There there is a maximal $v\in D\cap\Lambda$ such that $v\not\in S$, but $\chi^{v}$ appears in $f$. By Lemma˜3.2 and Lemma˜3.3, if $\chi^{v^{\prime}}$ appears in $f$ with $v^{\prime}\neq v$, then $v\not\in\operatorname{supp}\Psi(\chi^{v^{\prime}})$; therefore, $v\in\operatorname{supp}\Psi(f)$. It follows that $\Psi(f)$ is not contained in $\mathbb{Z}\left\langle\chi^{s}:s\in S\right\rangle$, hence is also not contained in
@@
Proof of Theorem˜1.2.
[…]
Proof of Corollary˜1.3.
Recall from [CLS11, Chapter 2] that an $M$-rational polytope $P\subset V$, gives rise to the projective toric variety $X_{P}:=\operatorname{Proj}\mathbb{Z}[S_{P}]$, where
$S_{P}:=\{(t,p):t\in\mathbb{Z}_{\geq 0},p\in tP\}\subset\mathbb{Z}\times M,$
graded by $\deg(t,p)=t$. Since $W$ preserves $M$ and $P$, it acts on $S_{P}$, hence on $\mathbb{Z}[S_{P}]$, yielding an action on $X_{P}$ by toric automorhpisms. The fundamental domain of the $W$-action on $\mathbb{R}\times V$ is $\mathbb{R}\times D$, so by Theorem˜1.2, there is an isomorphism of graded rings
$\mathbb{Z}[S_{D\cap P}]=\mathbb{Z}[(\mathbb{R}\times D)\cap S_{P}]\to\mathbb{Z}[S_{P}]^{W}.$
@@
For instance, if $u=e_{1}+3e_{2}=4z+2\lambda$, then $\Psi(\chi^{u})=\chi^{4z}\left(\chi^{\lambda}+\chi^{s_{\alpha}(\lambda)}\right)^{2}=\chi^{4z+2\lambda}+2\chi^{4z}+\chi^{4z-2\lambda}$. See Fig.˜1. Geometrically, we have that $\mathbb{A}^{2}/{S_{2}}\cong\mathbb{A}^{2}$.
[…]